A garden area is 30ft long and 20ft wide. A path of uniform width is set around the edge. If the remaining garden area is 400ft squared, what is the width of the path....

Answer by malakumar_kos@yahoo.com(315)


Area of the garden =30.20 = 600ft^2
area of the remaining garden = 400ft^2
Area of the path = 600-400=200ft^2
let the width of the path be =xft its length = x+10 ft
Area = l.b x(x+10)=200 x^2+10x-200=0 x^2+20x-10x-200=0
x(x+20)-10(x+20)=0
therefore (x+20)(x-10)=0 x-10=0 therefore x=10 and x+10=10+10=20

width of the path =10ft and its length =20ft